Part A MISSED THIS? Read Section 19.6 (Pages 863 - 867); Watch KCV 19.6, IWE 19.4. , T = 291 K Calculate the change in GibbsPart D T=397 K AHrxn= -90. kJ ASixnF 146 J/K Express your answer as an integer. Be a Ole ? AGIxn = Submit Request Answer
Answer #1

According to Gibbs Helmholtz equation,

∆G = ∆H - T∆S

1. ∆H ,= -90 kj. ∆S = 146 J/K. T = 291 K

1 Kj = 1000 J

∆H = -90000J

∆G = -90000 - 291×146

= -90000 - 42486

= -132486 J

= -132.5 KJ

∆G = -133 Kj

2. ∆G = -90000 - 851×(-146)

= -90000 + 124246

= 214246 J

= 241.3 KJ

∆G = 241 Kj

3. ∆G = 90000 - 291×(-146)

= 90000 + 42486

= 132486 j

= 132.5 kj

∆G = 133 Kj

4. ∆G = -90000 - 397 × (146)

= -90000 - 57962

= -147962 j

= -147.96 Kj

∆G = -148 kj